OpenStudy (dr.gray):

a wheel has a radius of 12 inches what is the circumference of the wheel?

OpenStudy (anonymous):

give me one sec

OpenStudy (welshfella):

2 * pi * r r = 12

OpenStudy (anonymous):

75.36

OpenStudy (anonymous):

u forgot to square it

OpenStudy (anonymous):

The diameter=24 Radius=12 2*pi*12

OpenStudy (welshfella):

just replace the r by 12 and work it out you can use 3.14 for pi

OpenStudy (anonymous):

75.36

OpenStudy (anonymous):

trust me

OpenStudy (welshfella):

= 2 * 12 * 3.14

OpenStudy (vijeya3):

well circumference of a circle=2*pi*r radius=12 inches put it in the formula above..... 75.36(approx)

OpenStudy (anonymous):

that is what i said

OpenStudy (welshfella):

milk333 plz dont give direct answers Read Code of Conduct.

OpenStudy (anonymous):

fan or medal me please

OpenStudy (anonymous):

and i did, and someone else did to